# Communication-Efficient Federated Optimization over Semi-Decentralized Networks

### v2

@@ A.5 Descent lemma
$\mathbb{E}[f(\overline{\boldsymbol{x}}^{k+1})]-\mathbb{E}[f(\overline{\boldsymbol{x}}^{k})]$ $\leq-\frac{\eta(T_{o}+1)}{4}\mathbb{E}[\|\nabla f(\overline{\boldsymbol{x}}^{k})\|_{2}^{2}]+\frac{10(T_{o}+1)L^{2}\eta}{n}\left(\mathbb{E}[\|\boldsymbol{\Phi}_{x}^{k}\|_{\mathsf{F}}^{2}]+(T_{o}+1)^{2}\eta_{l}^{2}\mathbb{E}[\|\boldsymbol{\Phi}_{y}^{k}\|_{\mathsf{F}}^{2}]\right)$
$+\left(64L^{2}T_{o}^{3}\eta\eta_{l}^{2}+\frac{3LT_{o}\eta^{2}}{2n}\right)\frac{\sigma^{2}}{b}+\frac{3L^{2}(T_{o}+1)^{3}\eta\eta_{l}^{2}}{n}\mathbb{E}[\|\overline{\boldsymbol{Y}}^{k}\|_{\mathsf{F}}^{2}].$ (29)
@@
We first handle the first term on the right-hand side. Let $\boldsymbol{g}^{k+1,t}=\frac{1}{n}\sum_{i=1}^{n}\boldsymbol{g}_{i}^{k+1,t}$. With (22), we have
@@
$=-\eta\sum_{t=0}^{T_{o}}\mathbb{E}[\nabla f(\overline{\boldsymbol{x}}^{k})^{\top}\left(\overline{\boldsymbol{g}}^{k+1,t}+\nabla f(\overline{\boldsymbol{x}}^{k})-\nabla f(\overline{\boldsymbol{x}}^{k})\right)]$
@@
$\leq\left(-\frac{\eta(T_{o}+1)}{2}+\frac{3LT_{o}^{2}\eta^{2}}{2}\right)\mathbb{E}[\|\nabla f(\overline{\boldsymbol{x}}^{k})\|_{2}^{2}]++\frac{\eta L^{2}}{2n}(1+3T_{o}L\eta)\sum_{t=0}^{T_{o}}\mathbb{E}[\|\boldsymbol{\Phi}_{x}^{k,t}\|_{\mathsf{F}}^{2}]+\frac{3LT_{o}\eta^{2}\sigma^{2}}{2nb}$
@@
A.6 Proof of Theorem 1
From the decent lemma, i.e., Lemma 7, summing (29) from $k=0$ to $k=K-1$ gives
$\frac{\eta(T_{o}+1)}{4}\sum_{k=0}^{K-1}\mathbb{E}[\|\nabla f(\overline{\boldsymbol{x}}^{k})\|_{2}^{2}]$ $\leq\tilde{f}+\frac{10(T_{o}+1)L^{2}\eta}{n}\left(\sum_{k=0}^{K-1}\mathbb{E}[\|\boldsymbol{\Phi}_{x}^{k}\|_{\mathsf{F}}^{2}]+(T_{o}+1)^{2}\eta_{l}^{2}\sum_{k=0}^{K-1}\mathbb{E}[\|\boldsymbol{\Phi}_{y}^{k}\|_{\mathsf{F}}^{2}]\right)$
$\quad+\left(64L^{2}T_{o}^{3}\eta\eta_{l}^{2}+\frac{3LT_{o}\eta^{2}}{2n}\right)\frac{K\sigma^{2}}{b}+\frac{3L^{2}(T_{o}+1)^{3}\eta\eta_{l}^{2}}{n}\sum_{k=0}^{K-1}\mathbb{E}[\|\overline{\boldsymbol{Y}}^{k}\|_{\mathsf{F}}^{2}],$ (33)
@@ A.6 Proof of Theorem 1
To show the convergence, we need to bound RHS of (33) by the following steps.
Step 1: formulating the dynamics of consensus errors and tracking errors.
For any $k>0$, let
@@
Assuming that $\eta\leq\frac{\lambda_{p}}{80L(T_{o}+1)}$ and $\eta_{l}\leq\frac{1}{8L(T_{o}+1)}$, we can formulate the dynamics by following Lemma 5 and Lemma 6,
@@
Step 2: showing the invertibility of $\boldsymbol{I}-(1-p)\boldsymbol{A}$.
To bound the sum of $\boldsymbol{\Phi}^{k}$, we have to control $\sum_{k=0}^{\infty}(1-p)^{k}\boldsymbol{A}^{k}$. To this end, the following lemma shows the spectral radius of $(1-p)\boldsymbol{A}$ is less than 1, implying that $\boldsymbol{I}-(1-p)\boldsymbol{A}$ is invertible.
@@
If $\eta\leq\frac{(1+p)\lambda_{p}^{2}}{80\sqrt{10}(T_{o}+1)L}$, it holds that $\rho\left((1-p)\boldsymbol{A}\right)<1$, where $\rho(\boldsymbol{A})$ represents the spectral radius of $\boldsymbol{A}$ defined in (36).
@@
Step 3: computing the cumulative consensus errors.
Because of the invertibility of $\boldsymbol{I}-(1-p)\boldsymbol{A}$, it follows [HJ12, Corollary 5.6.16] that
$\sum_{k=0}^{\infty}(1-p)^{k}\boldsymbol{A}^{k}=\left(\boldsymbol{I}-(1-p)\boldsymbol{A}\right)^{-1}$
@@
$\sum_{k=0}^{K}\boldsymbol{\Phi}^{k}\leq\left(\boldsymbol{I}-(1-p)\boldsymbol{A}\right)^{-1}\boldsymbol{\Phi}^{0}+\boldsymbol{C}\sum_{k=0}^{K-1}\boldsymbol{e}^{k}.$ (43)
@@
where $\det(\boldsymbol{A})$ means the determinant of $\boldsymbol{A}$ and $\operatorname{adj}(\boldsymbol{A})$ represents the adjugate of $\boldsymbol{A}$. If the step-size further satisfies $\eta\leq\frac{(1+p)\lambda_{p}^{2}}{360(T_{o}+1)L}$,
$\det(\boldsymbol{I}-(1-p)\boldsymbol{A})$ $=\left(1-\frac{(1-p)(1+(1+p)\lambda^{2})}{2}\right)^{2}-\frac{16000\lambda^{4}(T_{o}+1)^{2}L^{2}(1-p)^{2}\eta^{2}}{\lambda_{p}^{2}}$
$=\left(\frac{(1+p)\lambda_{p}}{2}\right)^{2}-\frac{16000\lambda^{4}(T_{o}+1)^{2}L^{2}(1-p)^{2}\eta^{2}}{\lambda_{p}^{2}}$
@@
Thus,
$(\boldsymbol{I}-(1-p)\boldsymbol{A})^{-1}$ $\leq\frac{8}{(1+p)^{2}\lambda_{p}^{2}}\begin{bmatrix}\frac{(1+p)\lambda_{p}}{2}&\frac{40(1-p)\lambda^{2}(T_{o}+1)^{2}\eta^{2}}{\lambda_{p}}\\ \frac{400(1-p)\lambda^{2}L^{2}}{\lambda_{p}}&\frac{(1+p)\lambda_{p}}{2}\end{bmatrix}$
$\leq\frac{4}{(1+p)^{2}\lambda_{p}^{3}}\begin{bmatrix}(1+p)\lambda_{p}^{2}&80(1-\lambda_{p})(T_{o}+1)^{2}\eta^{2}\\ 800(1-\lambda_{p})L^{2}&(1+p)\lambda_{p}^{2}\end{bmatrix}$
@@
$\boldsymbol{C}$ $\leq\frac{240(1-\lambda_{p})}{(1+p)^{2}\lambda_{p}^{4}}\begin{bmatrix}4c_{\lambda,1}L^{2}(T_{o}+1)^{4}\eta^{2}\eta_{l}^{2}&6nc_{\lambda,2}(T_{o}+1)^{2}\eta^{2}\\ 3c_{\lambda,3}L^{2}(T_{o}+1)^{2}\eta^{2}&3nc_{\lambda,4}\end{bmatrix}$
@@
$c_{\lambda,1}$ $=\left((1+p)\lambda_{p}^{2}+60(1-\lambda_{p})\eta_{c}^{2}\right),$
$c_{\lambda,2}$ $=\left((1+p)\lambda_{p}^{2}+40(1-\lambda_{p})\right),$
$c_{\lambda,3}$ $=\left((1+p)\lambda_{p}^{2}+1100(1-\lambda_{p})L^{2}(T_{o}+1)^{2}\eta_{l}^{2}\right),$
$c_{\lambda,4}$ $=\left((1+p)\lambda_{p}^{2}+1600(1-\lambda_{p})L^{2}(T_{o}+1)^{2}\eta^{2}\right).$
@@
$\quad+\frac{960c_{\lambda,1}(1-\lambda_{p})L^{2}(T_{o}+1)^{4}\eta^{2}\eta_{l}^{2}}{(1+p)^{2}\lambda_{p}^{4}}\sum_{k=0}^{K-1}\mathbb{E}[\|\overline{\boldsymbol{Y}}^{k}\|_{\mathsf{F}}^{2}]$
$\quad+\frac{1440c_{\lambda,2}(1-\lambda_{p})(T_{o}+1)^{2}\eta^{2}}{b(1+p)^{2}\lambda_{p}^{4}}nK\sigma^{2},$
@@
$\quad+\frac{720c_{\lambda,3}(1-\lambda_{p})L^{2}(T_{o}+1)^{2}\eta^{2}}{(1+p)^{2}\lambda_{p}^{4}}\sum_{k=0}^{K-1}\mathbb{E}[\|\overline{\boldsymbol{Y}}^{k}\|_{\mathsf{F}}^{2}]$
$\quad+\frac{720c_{\lambda,4}(1-\lambda_{p})}{b(1+p)^{2}\lambda_{p}^{4}}nK\sigma^{2}.$
Step 4: finishing up.
Substituting (A.6) into the (33), we have
$\frac{\eta(T_{o}+1)}{4}\sum_{k=0}^{K-1}\mathbb{E}[\|\nabla f(\overline{\boldsymbol{x}}^{k})\|_{2}^{2}]$
$\leq\tilde{f}+\frac{80c_{\lambda,1}L^{2}(T_{o}+1)^{3}\eta\eta_{l}^{2}}{n(1+p)^{2}\lambda_{p}^{3}}\mathbb{E}[\|\boldsymbol{\Phi}_{y}^{0}\|_{\mathsf{F}}^{2}]+\frac{3c_{2}L^{2}(T_{o}+1)^{3}\eta\eta_{l}^{2}}{n(1+p)^{2}\lambda_{p}^{4}}\sum_{k=0}^{K-1}\mathbb{E}[\|\overline{\boldsymbol{Y}}^{k}\|_{\mathsf{F}}^{2}]$
$\quad+\left(\underset{D_{1}}{\underbrace{\frac{7200c_{3}(1-\lambda_{p})L^{2}(T_{o}+1)^{3}\eta\eta_{l}^{2}}{(1+p)^{2}\lambda_{p}^{4}}+64L^{2}(T_{o}+1)^{3}\eta\eta_{l}^{2}}}+\frac{3LT_{o}\eta^{2}}{2n}\right)K\sigma^{2}/b,$ (45)
where
$c_{2}$ $=6400\max\{c_{\lambda,1},c_{\lambda,3}\}(1-\lambda_{p})L^{2}(T_{o}+1)^{2}\eta^{2}+(1+p)^{2}\lambda_{p}^{4}\leq 5(1+p)^{2}\lambda_{p}^{4}$
$c_{3}$ $=2c_{\lambda,2}\eta_{c}^{2}+c_{\lambda,4}\leq 84\eta_{c}^{2}+2(1+p)\lambda_{p}^{2}.$
Together with (15) and (44a), we have
@@
+($\quad+\underset{D_{2}}{\underbrace{\left(\frac{1440c_{\lambda,2}(1-\lambda_{p})L^{2}(T_{o}+1)^{2}\eta^{2}}{(1+p)^{2}\lambda_{p}^{4}}+\frac{1}{n}\right)}}6nK\sigma^{2}/b$
@@
$\frac{3\cdot 960c_{\lambda,1}L^{4}(T_{o}+1)^{4}\eta^{2}\eta_{l}^{2}}{(1+p)^{2}\lambda_{p}^{4}}\leq\frac{1}{2}.$
Because $c_{2}\leq 5(1+p)^{2}\lambda_{p}^{4}$, by further assuming that $\eta_{l}\leq\frac{1}{27L(T_{o}+1)}$ we have
$\frac{18c_{2}L^{2}(T_{o}+1)^{3}\eta\eta_{l}^{2}}{(1+p)^{2}\lambda_{p}^{4}}$ $\leq\frac{\eta(T_{o}+1)}{8},$
$\frac{5760c_{2}(1-p)\lambda^{2}L^{4}(T_{o}+1)^{5}\eta^{3}\eta_{l}^{2}}{n(1+p)^{4}\lambda_{p}^{7}}$ $\leq\alpha_{0}\frac{(1-p)L^{2}(T_{o}+1)^{3}\eta^{3}}{n(1+p)^{2}\lambda_{p}^{3}}=\frac{\alpha_{0}(1-p)(T_{o}+1)\eta}{n},$
$\frac{18c_{2}L^{2}(T_{o}+1)^{3}\eta\eta_{l}^{2}}{(1+p)^{2}\lambda_{p}^{4}}D_{2}$ $\leq\alpha_{1}L^{2}(T_{o}+1)^{3}\eta\eta_{l}^{2},$
for some positive absolute constant $\alpha_{0},\alpha_{1}$.
Assume that $\eta_{c}=\alpha\sqrt{1+p}\lambda_{p}$ and $\eta_{l}\leq\frac{\sqrt{1+p}\lambda_{p}}{360\alpha L(T_{o}+1)}$ for any positive $\alpha>1$. Then,
$\frac{80c_{\lambda,1}L^{2}(T_{o}+1)^{3}\eta\eta_{l}^{2}}{n(1+p)^{2}\lambda_{p}^{3}}$ $\leq\alpha_{2}\left(\frac{(T_{o}+1)\eta}{n}+\frac{(1-p)L^{2}(T_{o}+1)^{3}\eta^{3}}{n(1+p)^{2}\lambda_{p}^{3}}\right)=\alpha_{2}\left(\frac{(T_{o}+1)\eta}{n}\right)$
$D_{1}$ $\leq\alpha_{3}\left(\frac{(1-p)L^{2}(T_{o}+1)^{3}\eta^{3}}{(1+p)^{2}\lambda_{p}^{4}}+L^{2}(T_{o}+1)^{3}\eta\eta_{l}^{2}\right)$
for some positive absolute constant $\alpha_{2},\alpha_{3}$.
Therefore,
@@
$\leq\tilde{f}+O\left(\frac{(T_{o}+1)\eta}{n}\right)\mathbb{E}[\|\boldsymbol{\Phi}_{y}^{0}\|_{\mathsf{F}}^{2}]+O\left(\frac{(1-p)L^{2}(T_{o}+1)^{2}\eta^{2}}{(1+p)^{2}\lambda_{p}^{4}}+L^{2}(T_{o}+1)^{2}\eta_{l}^{2}+\frac{3L\eta}{2n}\right)\frac{K(T_{o}+1)\eta\sigma^{2}}{b},$
which is equivalent to
$\frac{1}{K}\sum_{k=0}^{K-1}\mathbb{E}[\|\nabla f(\overline{\boldsymbol{x}}^{k})\|_{2}^{2}]\leq O\left(\frac{\tilde{f}}{\eta T_{o}K}\right)+O\left(\frac{1}{nK}\right)\mathbb{E}[\|\boldsymbol{\Phi}_{y}^{0}\|_{\mathsf{F}}^{2}]+O\left(\frac{(1-p)L^{2}T_{o}^{2}\eta^{2}}{(1+p)^{2}\lambda_{p}^{4}}+L^{2}T_{o}^{2}\eta_{l}^{2}+\frac{L\eta}{n}\right)\frac{\sigma^{2}}{b}.$
Appendix B Proof of Corollary 1
Lemma 9 (Modified Lemma 17 in [${}^{+}$20]).
For any parameters $r_{0}\geq 0,a_{1}\geq 0,a_{2}\geq 0,a_{3}\geq 0$, if $K$ is sufficiently large s.t. $\eta=\min\left\{\left(\frac{r_{0}}{a_{1}K}\right)^{\frac{1}{2}},\left(\frac{r_{0}}{a_{2}K}\right)^{\frac{1}{3}},\left(\frac{r_{0}}{a_{3}K}\right)^{\frac{1}{3}}\right\}\leq\bar{\eta}$, i.e.,
$K\geq\max\left\{\frac{r_{0}}{a_{1}\bar{\eta}^{2}},\,\frac{r_{0}}{a_{2}\bar{\eta}^{3}},\,\frac{r_{0}}{a_{3}\bar{\eta}^{3}}\right\}$

### v4

@@ A.5 Descent lemma
$\mathbb{E}[f(\overline{\boldsymbol{x}}^{k+1})]-\mathbb{E}[f(\overline{\boldsymbol{x}}^{k})]$ $\leq-\frac{\eta(T_{o}+1)}{4}\mathbb{E}[\|\nabla f(\overline{\boldsymbol{x}}^{k})\|_{2}^{2}]+\frac{3L^{2}(T_{o}+1)^{3}\eta\eta_{l}^{2}}{n}\mathbb{E}[\|\overline{\boldsymbol{Y}}^{k}\|_{\mathsf{F}}^{2}]$
$\quad+\frac{10(T_{o}+1)L^{2}\eta}{n}\left(\mathbb{E}[\|\boldsymbol{\Phi}_{x}^{k}\|_{\mathsf{F}}^{2}]+(T_{o}+1)^{2}\eta_{l}^{2}\mathbb{E}[\|\boldsymbol{\Phi}_{y}^{k}\|_{\mathsf{F}}^{2}]\right)$
$\quad+\left(64L^{2}T_{o}^{3}\eta\eta_{l}^{2}+\frac{3LT_{o}\eta^{2}}{2n}\right)\frac{\sigma^{2}}{b},$ (25)
for any $k\geq 0$.
@@
We first handle the first term on the right-hand side. Let $\boldsymbol{g}^{k+1,t}=\frac{1}{n}\sum_{i=1}^{n}\boldsymbol{g}_{i}^{k+1,t}$. With Lemma 1 and the (9a) in hand, we have
@@
$\leq\left(-\frac{\eta(T_{o}+1)}{2}+\frac{3LT_{o}^{2}\eta^{2}}{2}\right)\mathbb{E}[\|\nabla f(\overline{\boldsymbol{x}}^{k})\|_{2}^{2}]+\frac{\eta L^{2}}{2n}(1+3T_{o}L\eta)\sum_{t=0}^{T_{o}}\mathbb{E}[\|\boldsymbol{\Phi}_{x}^{k,t}\|_{\mathsf{F}}^{2}]+\frac{3LT_{o}\eta^{2}\sigma^{2}}{2nb}$
@@
Appendix B Proof of Theorem 1
In this section, we present the convergence analysis of PISCO. From the descent lemma, i.e., Lemma 7, summing (25) from $k=0$ to $k=K-1$ gives
$\frac{\eta(T_{o}+1)}{4}\sum_{k=0}^{K-1}\mathbb{E}[\|\nabla f(\overline{\boldsymbol{x}}^{k})\|_{2}^{2}]$ $\leq\tilde{f}+\frac{3L^{2}(T_{o}+1)^{3}\eta\eta_{l}^{2}}{n}\sum_{k=0}^{K-1}\mathbb{E}[\|\overline{\boldsymbol{Y}}^{k}\|_{\mathsf{F}}^{2}]$
$+\frac{10(T_{o}+1)L^{2}\eta}{n}\sum_{k=0}^{K-1}\left(\mathbb{E}[\|\boldsymbol{\Phi}_{x}^{k}\|_{\mathsf{F}}^{2}]+(T_{o}+1)^{2}\eta_{l}^{2}\mathbb{E}[\|\boldsymbol{\Phi}_{y}^{k}\|_{\mathsf{F}}^{2}]\right)$
$+\left(64L^{2}T_{o}^{3}\eta\eta_{l}^{2}+\frac{3LT_{o}\eta^{2}}{2n}\right)\frac{K\sigma^{2}}{b},$ (29)
@@ Appendix B Proof of Theorem 1
To show the convergence of $\frac{1}{K}\sum_{k=0}^{K-1}\mathbb{E}[\|\nabla f(\overline{\boldsymbol{x}}^{k})\|_{2}^{2}]$, we need to bound right hand side of (29). To this end, we first formulate the dynamics of consensus errors and tracking errors. For any $k>0$, let
@@
Assuming that $\eta\leq\frac{\lambda_{p}}{80L(T_{o}+1)}$ and $\eta_{l}\leq\frac{1}{8L(T_{o}+1)}$, we can formulate the dynamics from Lemma 5 and Lemma 6,
@@
To control $\sum_{k=0}^{\infty}(1-p)^{k}\boldsymbol{A}^{k}$ in (33), we then establish the following lemma implying that $\boldsymbol{I}-(1-p)\boldsymbol{A}$ is invertible.
@@
If $\eta\leq\frac{(1+p)\lambda_{p}^{2}}{80\sqrt{10}(T_{o}+1)L}$,
$\rho\left((1-p)\boldsymbol{A}\right)<1,$
where $\rho(\boldsymbol{A})$ denotes the spectral radius of $\boldsymbol{A}$ defined in (30).
@@
Due to the invertibility of $\boldsymbol{I}-(1-p)\boldsymbol{A}$, it follows [HJ12, Corollary 5.6.16] that
$\sum_{k=0}^{\infty}(1-p)^{k}\boldsymbol{A}^{k}=\left(\boldsymbol{I}-(1-p)\boldsymbol{A}\right)^{-1},$
@@
$\sum_{k=0}^{K}\boldsymbol{\Phi}^{k}\leq\left(\boldsymbol{I}-(1-p)\boldsymbol{A}\right)^{-1}\boldsymbol{\Phi}^{0}+\boldsymbol{C}\sum_{k=0}^{K-1}\boldsymbol{e}^{k},$ (35)
@@
where $\det(\boldsymbol{A})$ means the determinant of $\boldsymbol{A}$ and $\operatorname{adj}(\boldsymbol{A})$ represents the adjugate of $\boldsymbol{A}$. If the step-size further satisfies $\eta\leq\frac{(1+p)\lambda_{p}^{2}}{360(T_{o}+1)L}$, we have
$\det(\boldsymbol{I}-(1-p)\boldsymbol{A})$ $=\left(\frac{(1+p)\lambda_{p}}{2}\right)^{2}-\frac{16000\lambda^{4}(T_{o}+1)^{2}L^{2}(1-p)^{2}\eta^{2}}{\lambda_{p}^{2}}$
@@
Then,
$(\boldsymbol{I}-(1-p)\boldsymbol{A})^{-1}$ $\leq\frac{4}{(1+p)^{2}\lambda_{p}^{3}}\begin{bmatrix}(1+p)\lambda_{p}^{2}&80(1-\lambda_{p})(T_{o}+1)^{2}\eta^{2}\\ 800(1-\lambda_{p})L^{2}&(1+p)\lambda_{p}^{2}\end{bmatrix},$
@@
$\boldsymbol{C}\!$ $\leq\!\frac{240(1-\lambda_{p})}{(1+p)^{2}\lambda_{p}^{4}}\begin{bmatrix}4c_{1}L^{2}(T_{o}+1)^{4}\eta^{2}\eta_{l}^{2}&6nc_{2}(T_{o}+1)^{2}\eta^{2}\\ 3c_{3}L^{2}(T_{o}+1)^{2}\eta^{2}&3nc_{4}\end{bmatrix},$
@@
$c_{1}$ $=\left((1+p)\lambda_{p}^{2}+60(1-\lambda_{p})\eta_{c}^{2}\right),$
$c_{2}$ $=\left((1+p)\lambda_{p}^{2}+40(1-\lambda_{p})\right),$
$c_{3}$ $=\left((1+p)\lambda_{p}^{2}+1100(1-\lambda_{p})L^{2}(T_{o}+1)^{2}\eta_{l}^{2}\right),$
$c_{4}$ $=\left((1+p)\lambda_{p}^{2}+1600(1-\lambda_{p})L^{2}(T_{o}+1)^{2}\eta^{2}\right).$
@@
$\quad+\frac{960c_{1}(1-\lambda_{p})L^{2}(T_{o}+1)^{4}\eta^{2}\eta_{l}^{2}}{(1+p)^{2}\lambda_{p}^{4}}\sum_{k=0}^{K-1}\mathbb{E}[\|\overline{\boldsymbol{Y}}^{k}\|_{\mathsf{F}}^{2}]$
$\quad+\frac{1440c_{2}(1-\lambda_{p})(T_{o}+1)^{2}\eta^{2}}{b(1+p)^{2}\lambda_{p}^{4}}nK\sigma^{2},$
and
@@
$\quad+\frac{720c_{3}(1-\lambda_{p})L^{2}(T_{o}+1)^{2}\eta^{2}}{(1+p)^{2}\lambda_{p}^{4}}\sum_{k=0}^{K-1}\mathbb{E}[\|\overline{\boldsymbol{Y}}^{k}\|_{\mathsf{F}}^{2}]$
$\quad+\frac{720c_{4}(1-\lambda_{p})}{b(1+p)^{2}\lambda_{p}^{4}}nK\sigma^{2}.$
Substituting (36) into the (29), we have
$\frac{\eta(T_{o}+1)}{4}\sum_{k=0}^{K-1}\mathbb{E}[\|\nabla f(\overline{\boldsymbol{x}}^{k})\|_{2}^{2}]$ $\leq\tilde{f}+\underset{T_{1}}{\underbrace{\frac{15L^{2}(T_{o}+1)^{3}\eta\eta_{l}^{2}}{n}\sum_{k=0}^{K-1}\mathbb{E}[\|\overline{\boldsymbol{Y}}^{k}\|_{\mathsf{F}}^{2}]}}+\underset{T_{2}}{\underbrace{\frac{80c_{1}L^{2}(T_{o}+1)^{3}\eta\eta_{l}^{2}}{n(1+p)^{2}\lambda_{p}^{3}}\mathbb{E}[\|\boldsymbol{\Phi}_{y}^{0}\|_{\mathsf{F}}^{2}]}}$ (37)
$\quad+\left(T_{3}+\frac{3LT_{o}\eta^{2}}{2n}\right)K\sigma^{2}/b$
where $T_{3}=\frac{7200(84\eta_{c}^{2}+2(1+p)\lambda_{p}^{2})(1-\lambda_{p})L^{2}(T_{o}+1)^{3}\eta\eta_{l}^{2}}{(1+p)^{2}\lambda_{p}^{4}}+64L^{2}(T_{o}+1)^{3}\eta\eta_{l}^{2}$. Together with (12) and (36a), we have
@@
$\quad+\left(\frac{1440c_{2}(1-\lambda_{p})L^{2}(T_{o}+1)^{2}\eta^{2}}{(1+p)^{2}\lambda_{p}^{4}}+\frac{1}{n}\right)6nK\sigma^{2}/b$
@@
$\frac{3\cdot 960c_{1}L^{4}(T_{o}+1)^{4}\eta^{2}\eta_{l}^{2}}{(1+p)^{2}\lambda_{p}^{4}}\leq\frac{1}{2}.$
By further assuming that $\eta_{l}\leq\frac{1}{27L(T_{o}+1)}$ we have
$T_{1}$ $\leq\frac{\eta(T_{o}+1)}{8}\sum_{k=0}^{K-1}\mathbb{E}[\|\nabla f(\overline{\boldsymbol{x}}^{k})\|_{2}^{2}]+\frac{\alpha_{0}(1-p)(T_{o}+1)\eta}{n}\mathbb{E}[\|\boldsymbol{\Phi}_{y}^{0}\|_{\mathsf{F}}^{2}]\!+\!\alpha_{1}L^{2}(T_{o}+1)^{3}\eta\eta_{l}^{2}K\frac{\sigma^{2}}{b},$
for some positive absolute constant $\alpha_{0},\alpha_{1}$. Suppose that $\eta_{c}=\alpha\sqrt{1+p}\lambda_{p}$ and $\eta_{l}\leq\frac{\sqrt{1+p}\lambda_{p}}{360\alpha L(T_{o}+1)}$ for some positive $\alpha>0.1$. Then,
$T_{2}\leq\alpha_{2}\left(\frac{(T_{o}+1)\eta}{n}\right)\mathbb{E}[\|\boldsymbol{\Phi}_{y}^{0}\|_{\mathsf{F}}^{2}],$
$T_{3}\leq\alpha_{3}\left(\frac{(1-p)L^{2}(T_{o}+1)^{3}\eta^{3}}{(1+p)^{2}\lambda_{p}^{4}}+L^{2}(T_{o}+1)^{3}\eta\eta_{l}^{2}\right)$
for some positive absolute constants $\alpha_{2}$ and $\alpha_{3}$. Therefore,
@@
$\leq\tilde{f}+\!O\!\left(\!\frac{(T_{o}+1)\eta}{n}\!\right)\!\mathbb{E}[\|\boldsymbol{\Phi}_{y}^{0}\|_{\mathsf{F}}^{2}]\!+\!O\!\left(\!\frac{(1-p)L^{2}(T_{o}+1)^{2}\eta^{2}}{(1+p)^{2}\lambda_{p}^{4}}\right.+\left.L^{2}(T_{o}+1)^{2}\eta_{l}^{2}+\frac{3L\eta}{2n}\right)\frac{K(T_{o}+1)\eta\sigma^{2}}{b},$
which is equivalent to (5).
Appendix C Proof of Corollary 1
By rearranging and relaxing (5), we have
$\frac{1}{K}\sum_{k=0}^{K-1}\mathbb{E}[\|\nabla f(\overline{\boldsymbol{x}}^{k})\|_{2}^{2}]$ $\leq O\left(\frac{\tilde{f}}{\eta_{l}\eta_{c}T_{o}K}+\frac{\eta_{l}\eta_{c}L\sigma^{2}}{nb}+\frac{1}{(1+p)\lambda_{p}^{2}}\frac{L^{2}T_{o}^{2}\eta_{l}^{2}\sigma^{2}}{b}\right.\left.+\frac{1}{nK}\mathbb{E}[\|\boldsymbol{\Phi}_{y}^{0}\|_{\mathsf{F}}^{2}]\right).$ (38)
To further fine-tune the step-size and obtain the exact convergence rate, we establish the following lemma, which is slightly different from the Lemma 17 in [KLB+20].
Lemma 9.
For any parameters $r_{0}\geq 0,a_{1}>0,a_{2}>0$, if $K$ is sufficiently large s.t. $\eta^{\prime}=\min\left\{\left(\frac{r_{0}}{a_{1}K}\right)^{\frac{1}{2}},\left(\frac{r_{0}}{a_{2}K}\right)^{\frac{1}{3}}\right\}\leq\bar{\eta}$, i.e.,
$K\geq\max\left\{\frac{r_{0}}{a_{1}\bar{\eta}^{2}},\frac{r_{0}}{a_{2}\bar{\eta}^{3}}\right\},$